OpenStudy (studytheworld):

Linear equation story problem: You are remodeling your kitchen. You've contacted two tiling companies who gladly told you how long it took their workers to tile a floor with a similar size. Jim completed half the floor in 7 hours. Pete completed the other half of the floor in 6 hours. If Pete can lay 15 more tiles per hour than Jim at what rate can Jim lay tiles? Let x represent Jim's rate. The answer is an equation.

OpenStudy (studytheworld):

@TheSmartOne Could you help?

OpenStudy (proheiper):

You still need help?

OpenStudy (studytheworld):

Yes please!

OpenStudy (proheiper):

Everyone learns at different paces, some problems seem simple to others, others find simple problems a little harder.

OpenStudy (proheiper):

I'm here to help, lets get started.

OpenStudy (proheiper):

This question is asking you to write an equation representing the rate of 2 people, they already got 1 person represented as a letter, can you tell me which one it is?

OpenStudy (studytheworld):

Jim

OpenStudy (proheiper):

what is representing jims rate?

OpenStudy (studytheworld):

x

OpenStudy (proheiper):

Ok so thats the first step of the equation.

OpenStudy (proheiper):

We need to let Pete represent something too, so lets use Pete as Y.

OpenStudy (studytheworld):

Okay

OpenStudy (proheiper):

Lets use \[y=x+15 \] as a starting equation to get to the final answer

OpenStudy (proheiper):

Y representing Pete, x representing Jim and 15 representing the "if" factor

OpenStudy (proheiper):

Do you understand?

OpenStudy (studytheworld):

Yes that makes sense.

OpenStudy (proheiper):

Ok next parts, we are almost done.

OpenStudy (proheiper):

The number of tiles Jim can lay in 7hrs is equal to the number of tiles Pete can lay in 6hrs.

OpenStudy (proheiper):

In an equation we could use that being said as 7x=6y, so if you use y=x+15 inside of 7x=6y you will get the answer.

OpenStudy (proheiper):

Do you want me to walk you through how to do it or do you know how to?

OpenStudy (studytheworld):

Hmm. So would it be like 7x=6(y+15) ?

OpenStudy (proheiper):

Correct!

OpenStudy (studytheworld):

Yay!!! Thanks so much for your help.